OpenStudy (anonymous):

A parallelogram is shown below: A parallelogram ABCD is shown with DC equal to 9 feet and the perpendicular distance between AB and DC equal to 8 over 9 foot. Part A: What is the area of the parallelogram? Show your work. (5 points) Part B: How can you decompose this parallelogram into two triangles? If this parallelogram was decomposed into two triangles, what would be the area of each triangle? (5 point @iGreen.

OpenStudy (anonymous):

@iGreen. ????????

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

In a Parallelogram OPPOSITE SIDES are EQUAL.

OpenStudy (anonymous):

Area = b*h where b = 9 feet and h = 8/9 foot.

OpenStudy (anonymous):

Hint : Divide by 9 i.e Area = b times h Area = 9 times 8/9

OpenStudy (anonymous):

9 gets cancelled.

OpenStudy (anonymous):

so whats the area you got?

OpenStudy (anonymous):

@sjman1 what area did you get for the parallelogram?

OpenStudy (anonymous):

@KenichiIvy i got 64 over 9

OpenStudy (anonymous):

How?

OpenStudy (anonymous):